OpenStudy (anonymous):

If F(x)=integral of sqrt(t^3+1) dt from 0 to x, then F'(2)=?

OpenStudy (anonymous):

\[F(x)=\int_0^x\sqrt{t^3+1}dt\\ f'(x)=\sqrt{x^3+1}{d\over dx}(x)=?\]

OpenStudy (anonymous):

this is differentiation under the integral scheme

terenzreignz (terenzreignz):

This is one of the rare instances that we can apply the first fundamental theorem of Calculus- near-directly. Allow me to state it for you... If \[\huge F(x) = \int_a^xf(t)dt\] Then \[\huge F'(x) = f(x)\]

terenzreignz (terenzreignz):

Of course, this assumes the continuity of f(t) over some interval containing [a,x] but that's a formality. Care to proceed, @Idealist ?

OpenStudy (anonymous):

The derivative I got is 3t^2/2sqrt(t^3+1) but do I do with that? Do I plug 2 into the derivative?

terenzreignz (terenzreignz):

No, see, \[\huge F(x)= \int_a^xf(t)dt =\int_0^x\sqrt{t^3 +1} \ dt \] So in this set up, what is f(t) ?

OpenStudy (anonymous):

I don't understand what you wrote in the formula, sorry.

terenzreignz (terenzreignz):

Okay, I hope you can see this correspondence... \[\huge F(x)= \int_a^x\color{red}{f(t)}dt =\int_0^x\color{red}{\sqrt{t^3 +1}} \ dt\]

OpenStudy (anonymous):

Wait a minute, so don't you just plug 2 into sqrt(t^3+1)?

terenzreignz (terenzreignz):

Yes, quite so :)

OpenStudy (anonymous):

Thank you. I'm sorry that I took a bit to understand.

terenzreignz (terenzreignz):

No worries :)